OpenStudy (anonymous):

Factor: 9x2 - 24x + 16 A. (3x - 4)2 B. (3x - 8)2 C. (3x - 8)(3x + 8) D. (3x - 4)(3x + 4)

OpenStudy (hoblos):

you've been asking the same question for hours!! you must know how to do that by yourself now can you try

OpenStudy (anonymous):

there is no answer in the choices

OpenStudy (anonymous):

it must be (3x-4)(3x+4)

OpenStudy (hoblos):

no it is not (3x-4)(3x+4).. it is one of the options but just let her try..

OpenStudy (anonymous):

If i knew how to do it , or understood i wouldn't be asking now would i ? NO

OpenStudy (anonymous):

aww (3x-4)(3x-4) hahah sorry

OpenStudy (anonymous):

ok

OpenStudy (hoblos):

you must use the formula a^2 -2ab + b^2 = (a-b)^2 do you know it??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

in my school of survival ... no one will going to trust to ... but you self is the one who trust on you

OpenStudy (hoblos):

(a-b)^2 = a^2 -2ab +b^2 (a+b)^2 = a^2 +2ab +b^2 (a-b)(a+b) = a^2 - b^2 these formulas you must know when factorizing